OpenStudy (anonymous):

Solve 3c + 4 = -20. -5.3 6 8 -8

OpenStudy (anonymous):

\(3c+4=-20\) \(3c=-20-4\) \(3c=-24\) \(c=\frac{-24}{3}=-8\)